The following table exhibits the income statement relationships

I. The following table exhibits the Income statement relationships. Fill in the appropriate blanks for each of the independent situations below.
Answer #1

Ɪ.

COMPANY A

a) Sales Revenue = Cost of goods sold + Gross profit

    Where Cost of goods sold = Beginning Inventory + Net Purchases - closing Inventory 

                                              =52,600+175,300-52,200

                                              = 175,700

Sales Revenue  = 175,700 + 95,300  = 271,000

 

b) Cost of Good sold = Beginning Inventory + Net Purchases - closing Inventory 

                                              =52,600+175,300-52,200

                                              = 175,700

 

c) Operating Expenses =  Gross profit - income before taxes

                                         =95,300 -6,000

                                          = 89,300

 

COMPANY B

 

d) Beginning Inventory = Cost of good sold -  Net Purchase + Closing Inventory

        Where Cost of goods sold = Sales - gross profit = 343,400 - 88,000  = 255,400

 

Beginning Inventory = 255,400- 255,600 +108,000 

                                 = 107,800

 

e) Cost of goods sold = Sales - gross profit

                                    = 343,400 - 88,000 

                                    = 255,400

 

f) Income Before taxes =   Gross profit - operating expenses

                                       =88,000 - 50,000

                                       = 38,000

 

COMPANY C

g) Net purchase =  Cost of goods sold - Opening inventory + Closing inventory

                             =387,000 -90,000 + 63,000

                               = 360,000

 

h) Gross Profit  = Sales - Cost of goods sold

                           =540,000 -387,000

                            = 153,000

 

i) Income Before taxes =  Gross Profit - operating expenses

                                        = 153,000 - 48,000

                                        = 105,000 


answered by: sidjn50